### v1

# Finite basis Kolmogorov-Arnold networks: domain decomposition for data-driven and physics-informed problems

Abstract
Kolmogorov-Arnold networks (KANs) have attracted attention recently as an alternative to multilayer perceptrons (MLPs) for scientific machine learning. However, KANs can be expensive to train, even for relatively small networks. Inspired by finite basis physics-informed neural networks (FBPINNs), in this work, we develop a domain decomposition method for KANs that allows for several small KANs to be trained in parallel to give accurate solutions for multiscale problems. We show that finite basis KANs (FBKANs) can provide accurate results with noisy data and for physics-informed training.
Keywords Kolmogorov-Arnold networks  $\cdot$ Physics-informed neural networks  $\cdot$ Domain decomposition
Figure 1: Graphical abstract.
1 Introduction
Scientific machine learning, predominantly developed using the theory of multilayer perceptrons (MLPs), has recently garnered a great deal of attention [1, 2, 3]. Kolmogorov-Arnold networks (KANs) [4] use the Kolmogorov-Arnold Theorem as inspiration for an alternative to MLPs. KANs offer advantages over MLPs in some cases, such as for continual learning and noisy data. While MLPs use trainable weights and biases on edges with fixed activation functions, KANs use trainable activation functions represented by splines. This switch may be beneficial because splines are easy to adjust locally and can switch between different resolutions. KANs have also been shown to provide increased interpretability over an MLP. Consequently, it is essential to provide a well-rounded analysis of the performance and applicability of KANs. An active area of research involves exploring how to utilize a KAN effectively and discerning in which cases a KAN might be preferred over an MLP. In this paper, we develop an architecture for KANs based on overlapping domain decompositions, which allows for more accurate solutions for complex problems.
Since the publication of [4], numerous variations of KANs have been released, including fractional KANs [5], deep operator KANs [6], physics-informed KANs (PI-KANs)[7], KAN-informed neural networks (KINNs)[8], temporal KANs [9], graph KANs [10, 11, 12], Chebyshev KANs (cKANs) [13], convolutional KANs [14], ReLU-KANs [15], among others. KANs have been applied to satellite image classification [16], abnormality detection [17], and computer vision [18]. While these results are promising, there remains much to learn about how KANs can best be utilized and optimized. Many of the recent variations employ alternative functions to the B-spline parametrization, while others extend KANs through well-developed methods from the field of machine learning, such as physics-informed neural networks (PINNs) [19], deep operator networks (DeepONets) [20], and recurrent neural networks. Because KANs still follow the general network architecture of an MLP, save for the switch between the edges and neurons, it is often straightforward to apply such techniques to KANs.
PINNs [19] have demonstrated success for a wide range of problems, including fluid dynamics [21, 22, 23, 24, 25], energy systems [26, 27, 28, 29, 30], and heat transfer [21]. PINNs allow for efficient and accurate solving of physical problems where robust data is lacking. This is accomplished by incorporating the governing differential equations of a system into the loss function of a neural network, which ensures that the resulting solution will satisfy the physical laws of the system. Although PINNs can successfully train with minimal data in many cases, there are notable cases in which challenges arise. For instance, PINNs can get stuck at the fixed points of a dynamical system while training [31]. PINNs are also difficult to train for multiscale problems, due to so-called spectral bias [32], or when the problem domain is large. Domain decomposition techniques can improve training in such cases [33, 34, 35, 36, 37]; for a general overview of domain decomposition approaches in scientific machine learning, see also [38, 39]. Rather than training the network over the entire domain, there are many approaches for breaking up the domain into subdomains and training networks on each subdomain. A popular choice for domain decomposition applied to PINNs is extended PINNs (XPINNs) introduced in [33], with a parallel version presented in [40], and extended in [34, 41]. XPINNs have different subnetworks for each domain, which are stitched together across boundaries and are trained in parallel. Another class of domain decomposition techniques for PINNs consists of time marching methods, where, in general, a network is initially trained for a small subdomain, and then the subdomain is slowly expanded until it contains the entire domain. Methods in this class include backward-compatible PINNs, which consist of a single network shared by all subdomains [35]. Also similar is the time marching approach in [36]. Finally, one particularly successful technique is finite basis PINNs (FBPINNs) [37, 42, 43, 44], which use partition of unity functions to weight the output of neural networks in each domain. In comparison to other techniques, FBPINNs offer simplicity because no boundary conditions are needed between adjacent domains. This work explores whether a similar technique can be used to enhance the training of a KAN for challenging cases.
In addition to domain decomposition, many methods have been developed to improve the training of PINNs, including adaptive weighting schemes [45, 32], residual-based attention [46], adaptive residual point selection [47, 48, 49, 50, 51, 52], causality techniques [53], multifidelity PINNs [54, 55], and hierarchical methods to learn more accurate solutions progressively [56, 57, 58, 59, 60, 61, 62]. A substantial advantage of the methods presented in this work is that these additional techniques could be applied in combination with FBKANs to improve the training further.
In this paper, we first introduce finite basis KANs (FBKANs) in section 2. In sections 3 and 4 we highlight some of the features of FBKANs applied to data-driven problems and physics-informed problems, respectively. We show that FBKANs can increase accuracy over KANs. One important feature of FBKANs is that the finite basis architecture serves as a wrapper around a KAN architecture. While we have chosen to focus on KANs as described in [4], most available extensions of KANs could be considered instead to increase accuracy or robustness further.
2 Methods
2.1 KANs
In [4], the authors proposed approximating a multivariate function $f(\mathbf{x})$ by a model of the form
$f(\mathbf{x})\approx\sum_{i_{n_{l}-1}=1}^{m_{n_{l}-1}}\varphi_{n_{l}-1,i_{n_{l}},i_{n_{l}-1}}\left(\sum_{i_{n_{l}-2}=1}^{m_{n_{l}-2}}\ldots\left(\sum_{i_{2}=1}^{m_{2}}\varphi_{2,i_{3},i_{2}}\left(\sum_{i_{1}=1}^{m_{1}}\varphi_{1,i_{2},i_{1}}\left(\sum_{i_{0}=1}^{m_{0}}\varphi_{0,i_{1},i_{0}}(x_{i_{0}})\right)\right)\right)\ldots\right),$ (1)
denoted as a Kolmogorov-Arnold network. Here, $n_{l}$ is the number of layers in the KAN, $\{m_{j}\}_{j=0}^{n_{l}}$ is the number of nodes per layer, and $\phi_{i,j,k}$ are the univariant activation functions. We denote the right-hand side of eq. 1 as $\mathcal{K}(x)$. The activation functions are polynomials of degree $k$ on a grid with $g$ grid points. They are represented by a weighted combination of a basis function $b(x)$ and a B-spline,
$\phi(x)=w_{b}b(x)+w_{s}\text{spline}(x)$ (2)
where
$b(x)=\frac{x}{1+e^{-x}}$
and
$\text{spline}(x)=\sum_{i}c_{i}B_{i}(x).$
Here, $B_{i}(x)$ is a polynomial of degree $k$. $c_{i},w_{b},$ and $w_{s}$ are trainable parameters.
KANs evaluate the B-splines on a precomputed grid. In one dimension, a domain $[a,b]$ a grid with $g_{1}$ intervals has grid points $\{t_{0}=a,t_{1},t_{2},\ldots,t_{g_{1}}=b\}$; cf. [4]. Grid extension [4] allows for fitting a new, fine-grained spline to a coarse-grained spline, increasing the expressivity of the KAN. The coarse splines are transferred to the fine splines following the procedure in [4].
In this work, we do not consider the method for enforcing sparsity as outlined in [4] and instead only consider the mean squared error in the loss function. Many variations of KANs have been proposed recently. However, in this work, we consider only the formulation as outlined in [4], although we note that the domain decomposition method presented here would work with many variants of KANs.
2.2 Physics-informed neural networks
A PINN is a neural network (NN) that is trained to approximate the solution of an initial-boundary value problem (IBVP) by optimizing loss terms accounting for initial conditions, boundary conditions, and a residual term using backpropagation to calculate derivatives. In particular, the residual term represents how well the output satisfies the governing differential equation [19]. More specifically, we aim at approximating the solution $f$ of a generic IBVP
$f_{t}+\mathcal{N}_{x}[f]$ $=0,$ $x\in\Omega,t\in[0,T],$ (3)
$f(x,t)$ $=g(x,t),$ $x\in\partial\Omega,t\in[0,T],$
$f(x,0)$ $=u(x),$ $x\in\Omega,$
over the open domain $\Omega\in\mathbb{R}^{d}$ with boundary $\partial\Omega.$ Here, $x$ and $t$ represent the spatial and temporal coordinates, respectively, $\mathcal{N}_{x}$ is the differential operator with respect to $x$, and $g$ and $u$ are given functions representing the boundary and initial conditions, respectively. The PINN model is then trained to minimize the mean squared error for the initial conditions, boundary conditions, and residual (physics) sampled at $N_{ic},N_{bc},N_{r}$ data points. We denote the sampling points for the initial conditions, boundary conditions, and residual as $\{x_{ic}^{i},u(x_{ic}^{i})\}_{i=1}^{N_{ic}}$, $\{(x_{bc}^{i},t_{bc}^{i}),g(x_{bc}^{i},t_{bc}^{i})\}_{i=1}^{N_{bc}}$, and $\{(x_{r}^{i},t_{r}^{i})\}_{i=1}^{N_{r}}$, respectively. We then optimize the following weighted loss function with respect to the trainable parameters $\theta$ of the NN model $f_{\theta}$:
$\mathcal{L}(\theta)$ $=\lambda_{ic}\mathcal{L}_{ic}(\theta)+\lambda_{bc}\mathcal{L}_{bc}(\theta)+\lambda_{r}\mathcal{L}_{r}(\theta),$ (4)
$\mathcal{L}_{ic}(\theta)$ $=\frac{1}{N_{ic}}\sum_{i=1}^{N_{ic}}\left(f_{\theta}(x_{ic}^{i},0)-u(x_{ic}^{i})\right)^{2},$ (5)
$\mathcal{L}_{bc}(\theta)$ $=\frac{1}{N_{bc}}\sum_{i=1}^{N_{bc}}\left(f_{\theta}(x_{bc}^{i},t_{bc}^{i})-g(x_{bc}^{i},t_{bc}^{i})\right)^{2},$ (6)
$\mathcal{L}_{r}(\theta)$ $=\frac{1}{N_{r}}\sum_{i=1}^{N_{r}}\left(\frac{\partial}{\partial t}f_{\theta}(x_{r}^{i},t_{r}^{i})+\mathcal{N}_{x}[f_{\theta}(x_{r}^{i},t_{r}^{i})]\right)^{2},$ (7)
where $\lambda_{ic},\lambda_{bc},\lambda_{r}$ are weights for the different loss terms. The choice of these weights can have a significant impact on the training process and model performance. While adaptive schemes are available for choosing the weights [45, 32, 63, 64], in this work, we set the weights manually.
2.3 FBPINNs
In FBPINNs [37, 42, 43, 44], we decompose the space domain $\Omega$ or the space-time domain $\Omega\times[0,T]$ into overlapping subdomains. Each overlapping subdomain $\Omega_{j}$ is the interior of the support of a corresponding function $\omega_{j}$, and all functions $\omega_{j}$ form a partition of unity. In particular, in the case of $L$ overlapping subdomains, we have
$\Omega=\bigcup_{j=1}^{L}\Omega_{j},\quad{\rm supp}(\omega_{j})=\overline{\Omega_{j}},\text{ and}\quad\sum_{j=1}^{L}\omega_{j}\equiv 1\text{ in }\Omega.$
In one dimension, a uniform overlapping domain decomposition with overlap ratio $\delta>1$ is given by subdomains
$\Omega_{j}=\left(\frac{(j-1)l-\delta l/2}{L-1},\frac{(j-1)l+\delta l/2}{L-1}\right),$
where $l=\max(x)-\min(x)$.
There are multiple ways to define the partition of unity functions. Here, we construct them based on the expression
$\omega_{j}=\frac{\hat{\omega}_{j}}{\sum_{j=1}^{L}\hat{\omega}_{j}},$ (8)
where
$\hat{\omega}_{j}(x)=\begin{cases}1&L=1,\\ \left[1+\cos\left(\pi(x-\mu_{j})/\sigma_{j}\right)\right]^{2}&L>1,\end{cases}$ (9)
with $\mu_{j}=l(j-1)/(L-1)$ and $\sigma_{j}=(\delta l/2)/(L-1)$ representing the center and half-width of each subdomain, respectively. An example of the one-dimensional partition of unity functions with $L=4$ is depicted in fig. 2.
Figure 2: Example the partition of unity functions on the domain $\Omega=[0,2]$ with $L=4$ subdomains.
In multiple dimensions, $\mathbf{x}\in\mathbb{R}^{d}$, we then obtain
$\hat{\omega}_{j}(x)=\begin{cases}1&L=1,\\ \Pi_{i=1}^{d}\left[1+\cos\left(\pi(x_{i}-\mu_{ij})/\sigma_{ij}\right)\right]^{2}&L>1,\end{cases}$ (10)
with $\mu_{ij}$ and $\sigma_{ij}$ representing the center and half-width of each subdomain $j$ in each direction $i$. An example of of a single partition of unity function with $d=2$ and $L=4$ is shown in fig. 3.
Figure 3: (Left) Example domain decomposition on the domain $\Omega=[-1,1]\times[-1,1]$ with $L=4$ subdomains. (Right) One example partition of unity function $\omega_{11}(x,y).$
Then, the FBPINN architecture reads
$f_{\theta}(x)=\sum_{j=1}^{L}\omega_{j}(x)f_{j,\theta^{j}}(x),$
where the $f_{j,\theta^{j}}$ is the neural network with parameters $\theta^{j}$ that corresponds to the subdomain $\Omega_{j}$; it is localized to $\Omega_{j}$ by multiplication with the partition of unity function $\omega_{j}$, which is zero outside of $\Omega_{j}$. The FBPINN model is trained in the same way as the PINN model, i.e., using initial, boundary, and residual loss terms; cf. section 2.2. Note that, in the original FBPINNs paper [37], hard enforcement of initial and boundary conditions is employed, such that only the residual loss term remains. Moreover, in [43], the approach has been extended to multilevel domain decompositions, and in [44], this domain decomposition approach has been applied to long-range time-dependent problems.
2.4 FBKANs
In FBKANs, we replace the NN used in FBPINNs with a KAN The function approximation in eq. 1 then becomes
$f(\mathbf{x})$ $\approx\sum_{j=1}^{L}\omega_{j}(x)\left[\sum_{i_{n_{l}-1}=1}^{m_{n_{l}-1}}\varphi^{j}_{n_{l}-1,i_{n_{l}},i_{n_{l}-1}}\left(\sum_{i_{n_{l}-2}=1}^{m_{n_{l}-2}}\ldots\left(\sum_{i_{1}=1}^{m_{1}}\varphi^{j}_{1,i_{2},i_{1}}\left(\sum_{i_{0}=1}^{m_{0}}\varphi^{j}_{0,i_{1},i_{0}}(x_{i_{0}})\right)\right)\ldots\right)\right]$ (11)
$=\sum_{j=1}^{L}\omega_{j}(x)\mathcal{K}^{j}(x;\theta^{j})$
where $\mathcal{K}^{j}(x;\theta^{j})$ denotes the $j^{th}$ KAN with trainable parameters $\theta^{j}$ and $\omega_{j}$ denotes the corresponding finite basis partition of unity function introduced in section 2.3. FBKANs are trained to minimize the loss function
$\mathcal{L}(\theta)=\lambda_{ic}\mathcal{L}_{ic}(\theta)+\lambda_{bc}\mathcal{L}_{bc}(\theta)+\lambda_{r}\mathcal{L}_{r}(\theta)+\lambda_{data}\mathcal{L}_{data}(\theta),$ (12)
composed of
$\mathcal{L}_{ic}(\theta)$ $=\frac{1}{N_{ic}}\sum_{i=1}^{N_{ic}}\left(\sum_{j=1}^{L}\omega_{j}(x)\mathcal{K}^{j}(x_{ic}^{i};\theta^{j})-u(x_{ic}^{i})\right)^{2},$
$\mathcal{L}_{bc}(\theta)$ $=\frac{1}{N_{bc}}\sum_{i=1}^{N_{bc}}\left(\sum_{j=1}^{L}\omega_{j}(x)\mathcal{K}^{j}(x_{bc}^{i};\theta^{j})-g(x_{bc}^{i},t_{bc}^{i})\right)^{2},$
$\mathcal{L}_{r}(\theta)$ $=\frac{1}{N_{r}}\sum_{i=1}^{N_{r}}\left(\frac{\partial}{\partial t}\left[\sum_{j=1}^{L}\omega_{j}(x)\mathcal{K}^{j}(x_{r}^{i};\theta^{j})\right]+\mathcal{N}_{x}\left[\sum_{j=1}^{L}\omega_{j}(x)\mathcal{K}^{j}(x_{r}^{i};\theta^{j})\right]\right)^{2},$
$\mathcal{L}_{data}(\theta)$ $=\frac{1}{N_{data}}\sum_{i=1}^{N_{data}}\left(\sum_{j=1}^{L}\omega_{j}(x)\mathcal{K}^{j}(x_{data}^{i};\theta^{j})-f(x_{data}^{i})\right)^{2}.$
Here, $\theta=\{\theta^{j}\}_{j=1}^{L}$ is the set of trainable parameters, and the loss function $\mathcal{L}$ contains both a term for data-driven training $\mathcal{L}_{data}$ and terms for physics-informed training $\mathcal{L}_{ic}$, $\mathcal{L}_{bc}$, and $\mathcal{L}_{r}$. In this way, FBKANs are adaptable to given problem characteristics.
For FBKANs, the grid of each local KAN $\mathcal{K}^{j}$ is set separately on the corresponding subdomain $\Omega^{j}$. To do so, we densely sample points $x\in\Omega$ and compute the partition of unity function $\omega_{j}(x)$. We then take the boundaries of each subdomain based on the partition of unity functions as
$a^{j}$ $=\min_{x\text{ s.t. }\omega^{j}(x)>0.0001}x,$
$b^{j}$ $=\max_{x\text{ s.t. }\omega^{j}(x)>0.0001}x.$
This gives a custom grid fit to each subdomain.
We evaluate the models using the relative $\ell_{2}$ error, calculated by
$\frac{||f(x)-\sum_{j=1}^{L}\omega_{j}(x)\mathcal{K}^{j}(x;\theta^{j})||_{2}}{||f(x)||_{2}}.$
Our implementation is based on Jax [65], using the Jax-KAN package [66] for the KAN implementation. For simplicity, we keep the hyperparameters of each KAN on each subdomain fixed, including the width and depth, reducing the total number of hyperparameters. However, in general, the parameters for each individual KAN $\mathcal{K}^{j}$ could be chosen differently.
3 Data-driven results
In this section, we consider data-driven problems, which are trained without physics meaning that $\lambda_{ic}=\lambda_{bc}=\lambda_{r}=0$. These results specifically illustrate how FBKANs can improve predictions for noisy data and multiscale oscillations.
3.1 Test 1
In this section, we consider a toy problem
$f(x)=\exp[\sin(0.3\pi x^{2})],$ (13)
on $x\in[0,8]$. This problem is designed to test both the scaling of FBKANs as we increase the number of subdomains, as well as the impact of noisy data on training FBKANs.
3.1.1 Scaling of FBKANs
[…]
Figure 4: Results with $N=2,8,$ and $32$ subdomains for eq. 13. (a) Training data and plot of exact $f(x)$. (b) Loss curves (eq. 12). (c) Plot of the outputs $f(x)$ and $\sum_{j=1}^{L}\omega_{j}(x)\mathcal{K}^{j}(x)$. (d) Pointwise errors $f(x)-\sum_{j=1}^{L}\omega_{j}(x)\mathcal{K}^{j}(x)$. (e) Scaling results for Test 1 with $L$ subdomains.
3.1.2 FBKANs with noisy data
One strength of KANs over MLPs is their increased accuracy for noisy data [4]. We now test FBKANs on noisy training data using four subdomains ($L=4$). Therefore, we sample 600 training points from a uniform distribution in $[0,8]$ and evaluate $f(x)$. Then, we add Gaussian white noise with zero mean and varying magnitude to $f(x)$, up to an average relative magnitude of $20\%$. This becomes the training set. The method is then tested on 1 000 evenly spaced points in $[0,8]$ without noise.
The training set and results are shown in fig. 5. In the noisiest case, with relative noise of 18.1% added to the training set, the KAN yields a relative $\ell_{2}$ error of 0.1404, whereas the FBKAN yields a relative error of 0.0646. For all noise levels tested, the FBKAN consistently has a lower relative error than the plain KAN in fig. 5, and the predictions are robust to noisy data.
Figure 5: Results for Test 1 with noisy training data; cf. eq. 13. (a) Example training data and plot of exact $f(x)$ with $9.6\%$ mean relative noise. (b) Loss curves (eq. 12) for an example training with $9.6\%$ mean relative noise. (c) Plot of the outputs $f(x)$ and $\sum_{j=1}^{L}\omega_{j}(x)\mathcal{K}^{j}(x)$ with $9.6\%$ mean relative noise. (d) Pointwise errors $f(x)-\sum_{j=1}^{L}\omega_{j}(x)\mathcal{K}^{j}(x)$ with $9.6\%$ mean relative noise. (e) Relative $\ell_{2}$ error of the KANs and FBKANs with respect to the magnitude of the noise added to the training data.
3.2 Test 2
We now consider the following equation
$f(x,y)=\sin(6\pi x^{2})\sin(8\pi y^{2}),$ (14)
for $(x,y)\in[0,1]\times[0,1].$ This data-driven example exhibits fine-scale oscillations. We test two cases: (1) a KAN and FBKAN with a fixed grid, denoted by KAN-1/FBKAN-1 and (2) a KAN and FBKAN with grid extension, denoted by KAN-2/FBKAN-2. The FBKAN has $L=4$ subdomains. The training set is composed of 10 000 points randomly sampled from a uniform distribution.
In both cases, we begin training with $g=5$ grid points. In the grid extension case, the grid increases every 600 iterations as listed in table 4, and the learning rate drops by $20\%$ each time the grid is updated. As can be observed in fig. 6, the final training loss for FBKAN-1 with a fixed grid is approximately the same as the training loss for KAN-2 with the grid extension approach, even though each FBKAN network has six times fewer grid points. This is also reflected in the relative errors reported in table 1. Comparing the KAN-1 and FBKAN-1 models in fig. 7, KAN-1 struggles to capture the fine-scale features of the data accuratelyand has a larger relative error. As can be seen in fig. 8, KAN-2 and FBKAN-2 both outperform their counterparts with a static grid, but FBKAN-2 is better able to capture the data than KAN-2.
Figure 6: Training loss curves for Test 2.
Name Grid type Relative error
KAN-1 Fixed grid $2.36\times 10^{-1}$
FBKAN-1 Fixed grid $7.43\times 10^{-2}$
KAN-2 Grid extension $8.10\times 10^{-2}$
FBKAN-2 Grid extension $2.27\times 10^{-2}$
Table 1: Relative $\ell_{2}$ errors for physics-informed Test 2.
Figure 7: Reference solution (left), predicted solutions (middle) and relative errors (right) for KAN-1 and FBKAN-1 for Test 2. Note that the error plots have different color scales.
Figure 8: Reference solution (left), predicted solutions (middle) and relative errors (right) for KAN-2 and FBKAN-2 for Test 2. Note that the error plots have different color scales.
4 Physics-informed results
4.1 Test 3
As the first physics-informed example, we consider the following one-dimensional problem with multiscale features:
$\frac{df}{dx}$ $=4\cos(4x)+40\cos(40x),$
$f(0)$ $=0,$
on $x\in[-4,4]$ with exact solution
$f(x)=\sin(4x)+\sin(40x).$ (15)
We sample 400 collocation points from a uniform distribution ($N_{r}=400$) at every training iteration. The grid starts with $g=5$ and increases by $5$ every 1 000 iterations. The KAN reaches a relative $\ell_{2}$ error of 0.2407 with $g_{final}=20$, whereas the FBKAN for four subdomains ($L=4$) reaches a relative $\ell_{2}$ error of 0.0898 for $g_{final}=20$. For eight subdomains ($L=8$), the FBKAN reaches a relative error of 0.0369 with $g_{final}=15$ and 0.0662 with $g_{final}=10$.
Figure 9: Results for Test 3 with $L=8$. The FBKAN clearly outperforms the KAN, reaching a lower loss with fewer grid points.
4.2 Test 4
We now test the Helmholtz equation
$\frac{\partial^{2}f}{\partial y^{2}}+\frac{\partial^{2}f}{\partial x^{2}}+k_{h}^{2}f-q(x,y)$ $=0,\ (x,y)\in[-1,1]\times[-1,1],$
$f(-1,y)=f(1,y)$ $=0,\ y\in[-1,1],$
$f(x,-1)=f(x,1)$ $=0,\ x\in[-1,1],$
with
$q(x,y)=-(a_{1}\pi)^{2}\sin(a_{1}\pi x)\sin(a_{2}\pi y)-(a_{2}\pi)^{2}\sin(a_{1}\pi x)\sin(a_{2}\pi y)+k_{h}\sin(a_{1}\pi x)\sin(a_{2}\pi y).$
In our tests, we vary $a_{1}$ and $a_{2}$.
For each choice of $(a_{1},a_{2})$ we consider three training cases. In the first case, we use higher-order splines with $k=5$ and a fixed grid with $g=5$ (denoted by KAN-1/FBKAN-1). In the second case, we take $k=3$ and use grid extension to increase $g$ by $5$ every $600$ iterations, starting with $g=5$ (denoted by KAN-2/FBKAN-2). We use one hidden layer with width 10 for both cases. Additionally, we test a third case, denoted by KAN-3/FBKAN-3, where we consider the same hyperparameters as KAN-1/FBKAN-1 but with width 5 in the hidden layer. The relative $\ell_{2}$ errors are reported in table 2. In all cases, the FBKAN outperforms the corresponding KAN. The grid extension KAN cases, where we increase the grid size over the training, comes at the expense of increasing the computational time by a factor of 1.4-1.5 in our tests. For $a_{1}=1,a_{2}=4$, the smaller FBKAN-3 network is sufficient to represent the solution accurately. However, for larger values of $a_{1}$ and $a_{2}$, the larger network in FBKAN-1 is necessary. The results for KAN-1/FBKAN-1 are shown in figs. 10 and 11 for $a_{1}=1,a_{2}=4$ and $a_{1}=4,a_{2}=4$, respectively. In fig. 12, we consider $a_{1}=a_{2}=6$ with $L=4,9$ and $16$, demonstrating the further refinement possible with additional subdomains.
$a_{1}=1$, $a_{2}=4$ $a_{1}=4$, $a_{2}=4$ $a_{1}=6$, $a_{2}=6$
KAN-1 0.0259 0.5465 1.1254
FBKAN-1, L=4 0.0102 0.0267 0.1151
FBKAN-1, L=9 0.0213 0.0239 0.0399
FBKAN-1, L=16 0.0037 0.0128 0.0321
KAN-2 0.0180 0.2045 0.5854
FBKAN-2 0.0112 0.0427 0.2272
KAN-3 0.3771 0.5488 1.2825
FBKAN-3 0.0214 0.2760 0.9797
Table 2: Relative $\ell_{2}$ errors for physics-informed Test 3.
Figure 10: Results for Test 4 with KAN-1 and FBKAN-1, and $a_{1}=1,a_{2}=4$, $L=4$. The solution along the line $x=0$ is given in the bottom left subfigure. Note that the error plots have different color scales.
Figure 11: Results for Test 4 with KAN-1 and FBKAN-1, and $a_{1}=4,a_{2}=4$, $L=4$. The solution along the line $x=0$ is given in the bottom left subfigure. Note that the error plots have different color scales.
Figure 12: Results for Test 4 with KAN-1 and FBKAN-1 and $a_{1}=a_{2}=6$, with $L=4$ and $L=16$.The solution along the line $x=0$ is given in the middle left subfigure. Note that the error plots have different color scales.
4.3 Test 5
Finally, we consider the wave equation
$\frac{\partial^{2}f}{\partial t^{2}}-c^{2}\frac{\partial^{2}f}{\partial x^{2}}$ $=0,\ (x,t)\in[0,1]\times[0,1],$
$f(0,t)$ $=0,\ t\in[0,1],$
$f(1,t)$ $=0,\ t\in[0,1],$
$f(x,0)$ $=\sin(\pi x)+0.5\sin(4\pi x),\ x\in[0,1],$
$f_{t}(x,0)$ $=0,\ x\in[0,1],$
which has the exact solution
$f(x,t)=\sin(\pi x)\cos(c\pi t)+0.5\sin(4\pi x)\cos(4c\pi t).$
We first consider $c=\sqrt{2}$. The KAN has a relative $\ell_{2}$ error of 0.1402, and the FBKAN with $L=4$ has a relative $\ell_{2}$ error of 0.0153, as illustrated in fig. 13. We then consider the harder case with $c=2$, shown in fig. 14. The KAN has a relative $\ell_{2}$ error of 0.1778 and the FBKAN with $L=4$ has a relative $\ell_{2}$ error of 0.0587.
Figure 13: Results for Test 5 with $c=\sqrt{2}$ and $L=4$. Note that the error plots have different color scales.
Figure 14: Results for Test 5 with $c=2$ and $L=4.$ Note that the error plots have different color scales.
5 Conclusions
We have developed domain decomposition-based KAN models for data-driven and physics-informed training with KANs; in accordance with FBPINNs, we denote them as FBKANs. FBKANs are scalable to complex problems and have a strong advantage over other domain decomposition-based approaches in that they do not require enforcement of transmission conditions between the subdomains via the loss function. They allow accurate training using an ensemble of small KANs combined using partition of unity functions, instead of a single large network.
One advantage of FBKANs is that they can be combined with existing techniques to improve the training of KANs and PI-KANs, including residual-based attention weights as introduced in [7], cKANs [13], deep operator KANs [6], and others. These methods would be directly compatible with the FBKAN framework. In future work, we will further examine the scalability of FBKANs and consider their application to higher dimensional problems. We will also consider multilevel FBKANs, following the multilevel FBPINNs [43] approach. Multilevel FBPINNs show improvement over FBPINNs for a large number of subdomains by providing a mechanism for global communication between subdomains. Multilevel FBKANs could offer a similar advantage, allowing robust training with an increasing number of subdomains.
6 Code and data availability
All code, trained models, and data required to replicate the examples presented in this paper will be released upon publication. Meanwhile, we have released code and Google Colab tutorials for FBKANs in Neuromancer [67] at https://github.com/pnnl/neuromancer/tree/feature/fbkans/examples/KANs, for the reader to explore the ideas implemented in this work.
7 Acknowledgements
The KAN diagrams in fig. 1 were developed with pykan [4]. This project was completed with support from the U.S. Department of Energy, Advanced Scientific Computing Research program, under the Scalable, Efficient and Accelerated Causal Reasoning Operators, Graphs and Spikes for Earth and Embedded Systems (SEA-CROGS) project (Project No. 80278) and under the Uncertainty Quantification for Multifidelity Operator Learning (MOLUcQ) project (Project No. 81739). The computational work was performed using PNNL Institutional Computing at Pacific Northwest National Laboratory. Pacific Northwest National Laboratory (PNNL) is a multi-program national laboratory operated for the U.S. Department of Energy (DOE) by Battelle Memorial Institute under Contract No. DE-AC05-76RL01830.
Appendix A Training parameters
All results in this paper are implemented in JAX [65] using the Jax-KAN [66] KAN implementation. All networks are trained with the ADAM optimizer. For all FBKANs we take the domain overlap $\delta=1.9$.
A.1 Test 1
parameter section 3.1.1 section 3.1.2
KAN architecture [1, 5, 1] [1, 5, 1]
$L$ 4 - 32 4
$g$ 5 5
$k$ 3 3
Learning rate 0.04 0.04
Iterations 4000 4000
$N_{data}$ 1200 600
Table 3: Hyperparameters used for the results in section 3.1.
A.2 Test 2
Parameter KAN-1/FBKAN-1 KAN-2/FBKAN-2
KAN architecture [2, 10, 1] [2, 5, 1]
$L$ 4 4
$g$ 5 [5, 10, 25, 30]
$g$ schedule - [0, 600, 1200, 1800]
$k$ 3 3
Initial learning rate 0.02 0.02
Learning rate scale - 0.8
Iterations 2400 2400
$N_{data}$ 10000 10000
Table 4: Hyperparameters used for the results in section 3.2. The grid ($g$) schedule denotes the iterations at which the grid is updated. The learning rate scale denotes the change to the learning rate at each grid update. KAN-1/FBKAN-1 use fixed grids.
A.3 Test 3
Parameter
KAN architecture [2, 10, 1]
$L$ 4, 8
$g$ [5, 10, 15, 20]
$g$ schedule [0, 1000, 2000, 3000]
$k$ 3
Initial learning rate 0.01
Learning rate scale 0.8
Iterations 4000
$N_{r}$ 400
$N_{ic}$ 1
$\lambda_{r}$ 1/40
$\lambda_{ic}$ 1
Table 5: Hyperparameters used for the results in section 4.1. The grid ($g$) schedule denotes the iterations at which the grid is updated. The learning rate scale denotes the change to the learning rate at each grid update.
A.4 Test 4
Parameter KAN-1/FBKAN-1 KAN-2/FBKAN-2 KAN-3/FBKAN-3
KAN architecture [2, 10, 1] [2, 10, 1] [2, 5, 1]
$L$ 4-16 4 4
$g$ 5 [5, 10, 15] 5
$g$ schedule - [0, 3000, 6000] -
$k$ 5 3 5
Initial learning rate 0.005 0.005 0.005
Learning rate scale - 0.8 -
Iterations $a_{1}=1,a_{2}=4$: 10000 $a_{1}=4,a_{2}=4$: 10000 $a_{1}=6,a_{2}=6$: 30000 10000 10000
$N_{r}$ 800 800 800
$N_{bc}$ 400 400 400
$\lambda_{r}$ 0.01 0.01 0.01
$\lambda_{ic}$ 1 1 1
Table 6: Hyperparameters used for the results in section 4.2. The grid ($g$) schedule denotes the iterations at which the grid is updated. The learning rate scale denotes the change to the learning rate at each grid update.
A.5 Test 5
Parameter $c=\sqrt{2}$ $c=2$
KAN architecture [2, 10, 1] [2, 10, 10, 1]
$L$ 4 4
$g$ 10 10
$k$ 5 5
Initial learning rate 0.001 0.0005
Iterations 60000 120000
$N_{r}$ 1000 1200
$N_{ic}$ 100 100
$N_{bc}$ 200 200
$\lambda_{r}$ 0.01 0.01
$\lambda_{ic}$ 1 1
$\lambda_{bc}$ 1 1
Table 7: Hyperparameters used for the results in section 4.3.

### v2

@@
 However, KANs can be expensive to train, even for relatively small networks.
-Inspired by finite basis physics-informed neural networks (FBPINNs), in this work, we develop a domain decomposition method for KANs that allows for several small KANs to be trained in parallel to give accurate solutions for multiscale problems.
-We show that finite basis KANs (FBKANs) can provide accurate results with noisy data and for physics-informed training.
-Keywords Kolmogorov-Arnold networks  $\cdot$ Physics-informed neural networks  $\cdot$ Domain decomposition
-Figure 1: Graphical abstract.
+Inspired by finite basis physics-informed neural networks (FBPINNs), in this work, we develop a domain decomposition-based architecture for KANs that allows for several small KANs to be trained in parallel to give accurate solutions for multiscale problems.
+We denote this new approach as finite basis KANs (FBKANs) and show that it yields accurate results for noisy data as well as physics-informed training.
+Keywords Kolmogorov-Arnold networks $\cdot$ Physics-informed neural networks $\cdot$ Domain decomposition $\cdot$ Network architecture $\cdot$ Multiscale problems $\cdot$ Finite basis PINNs
+Figure 1: Graphical abstract: The computational domain is decomposed into overlapping subdomains, and an individual KAN model is defined on each subdomain.
+Then, the global model output is obtained by scaling the outputs of the local KAN models with partition of unity functions and summing up the local contributions.
 1 Introduction
-Scientific machine learning, predominantly developed using the theory of multilayer perceptrons (MLPs), has recently garnered a great deal of attention [1, 2, 3].
-Kolmogorov-Arnold networks (KANs) [4] use the Kolmogorov-Arnold Theorem as inspiration for an alternative to MLPs.
-KANs offer advantages over MLPs in some cases, such as for continual learning and noisy data.
-While MLPs use trainable weights and biases on edges with fixed activation functions, KANs use trainable activation functions represented by splines.
-This switch may be beneficial because splines are easy to adjust locally and can switch between different resolutions.
+Scientific machine learning, predominantly developed using the theory of multilayer perceptrons (MLPs), has garnered a great deal of attention [karniadakis2021physics, baker2019workshop, carter2023advanced] over recent years.
+An alternative model architecture, which is based on the Kolmogorov-Arnold representation theorem [kolmogorov1957representations], has recently been introduced in [liu2024kan].
+KANs offer advantages over MLPs in some contexts, such as for continual learning and noisy data.
+While MLPs use trainable weights and biases on edges with fixed activation functions, KANs employ trainable activation functions represented by splines.
+This switch may be beneficial because splines are easy to adjust locally and adapt to different resolutions.
 KANs have also been shown to provide increased interpretability over an MLP.
-Consequently, it is essential to provide a well-rounded analysis of the performance and applicability of KANs.
-An active area of research involves exploring how to utilize a KAN effectively and discerning in which cases a KAN might be preferred over an MLP.
-In this paper, we develop an architecture for KANs based on overlapping domain decompositions, which allows for more accurate solutions for complex problems.
-Since the publication of [4], numerous variations of KANs have been released, including fractional KANs [5], deep operator KANs [6], physics-informed KANs (PI-KANs)[7], KAN-informed neural networks (KINNs)[8], temporal KANs [9], graph KANs [10, 11, 12], Chebyshev KANs (cKANs) [13], convolutional KANs [14], ReLU-KANs [15], among others.
-KANs have been applied to satellite image classification [16], abnormality detection [17], and computer vision [18].
-While these results are promising, there remains much to learn about how KANs can best be utilized and optimized.
-Many of the recent variations employ alternative functions to the B-spline parametrization, while others extend KANs through well-developed methods from the field of machine learning, such as physics-informed neural networks (PINNs) [19], deep operator networks (DeepONets) [20], and recurrent neural networks.
-Because KANs still follow the general network architecture of an MLP, save for the switch between the edges and neurons, it is often straightforward to apply such techniques to KANs.
-PINNs [19] have demonstrated success for a wide range of problems, including fluid dynamics [21, 22, 23, 24, 25], energy systems [26, 27, 28, 29, 30], and heat transfer [21].
-PINNs allow for efficient and accurate solving of physical problems where robust data is lacking.
+Consequently, it is essential to provide a thorough analysis of the performance and applicability of KANs.
+An active area of research involves exploring how to utilize a KAN effectively and discerning in which cases KANs might be preferred over classical MLPs.
+In this paper, we develop an architecture for KANs based on overlapping domain decompositions, which allows for more accurate solutions for complex, and particularly multiscale, problems.
+Since the publication of [liu2024kan], numerous variations of KANs have been released, including fractional KANs [aghaei2025fkan], deep operator KANs [abueidda2025deepokan], physics-informed KANs (PI-KANs)[shukla2024comprehensive], KAN-informed neural networks (KINNs)[wang2025kolmogorov], temporal KANs [genet2024tkan], graph KANs [kiamari2024gkan, decarlo2024kolmogorovarnoldgraphneuralnetworks, bresson2024kagnnskolmogorovarnoldnetworksmeet], Chebyshev KANs (cKANs) [ss2024chebyshev], convolutional KANs [Bodner2024], ReLU-KANs [qiu2024relu], among others.
+KANs have been applied to satellite image classification [cheon2024kolmogorov], abnormality detection [huang2024abnormality], and computer vision [azam2024suitability].
+Although these results are promising, many open questions remain about how KANs can be best utilized and optimized.
+Many of the recent variations employ alternative functions to the B-spline parametrization, while others extend KANs through well-developed methods from the field of machine learning, such as physics-informed neural networks (PINNs) [raissi2019physics], deep operator networks (DeepONets) [lu2019deeponet], and recurrent neural networks.
+Because KANs still follow the general network architecture of an MLP, aside from the switch between the edges and neurons, it is often straightforward to apply such techniques to KANs.
+PINNs [raissi2019physics] have demonstrated success for a wide range of problems, including fluid dynamics [cai2021physics, jin2021nsfnets, raissi2020hidden, mahmoudabadbozchelou2022nn, almajid2022prediction], energy systems [chen2023physics, misyris2020physics, huang2022applications, moya2023dae, bento2023physics], and heat transfer [cai2021physics].
+PINNs allow for the efficient and accurate solution of physical problems where robust data is lacking.
 This is accomplished by incorporating the governing differential equations of a system into the loss function of a neural network, which ensures that the resulting solution will satisfy the physical laws of the system.
-Although PINNs can successfully train with minimal data in many cases, there are notable cases in which challenges arise.
-For instance, PINNs can get stuck at the fixed points of a dynamical system while training [31].
-PINNs are also difficult to train for multiscale problems, due to so-called spectral bias [32], or when the problem domain is large.
-Domain decomposition techniques can improve training in such cases [33, 34, 35, 36, 37]; for a general overview of domain decomposition approaches in scientific machine learning, see also [38, 39].
+Although PINNs can successfully train with little or no data in many cases, there are notable cases in which challenges arise.
+For instance, PINNs can get stuck at the fixed points of dynamical systems during training [rohrhofer2023ds].
+PINNs are also difficult to train for multiscale problems, due to the so-called spectral bias [rahaman_spectral_2019, wang2022and], or when the problem domain is large.
+Domain decomposition techniques can improve training in such cases [jagtap2020extended, penwarden2023unified, mattey2022novel, wight2020solving, moseley2023finite]; for a general overview of domain decomposition approaches in scientific machine learning, see also [heinlein_combining_2021, klawonn_machine_2024].
 Rather than training the network over the entire domain, there are many approaches for breaking up the domain into subdomains and training networks on each subdomain.
-A popular choice for domain decomposition applied to PINNs is extended PINNs (XPINNs) introduced in [33], with a parallel version presented in [40], and extended in [34, 41].
-XPINNs have different subnetworks for each domain, which are stitched together across boundaries and are trained in parallel.
+A popular choice for domain decomposition applied to PINNs is extended PINNs (XPINNs) introduced in [jagtap2020extended], with a parallel version presented in [shukla2021parallel], and extended in [penwarden2023unified, hu2022extended].
+XPINNs employ a separate subnetwork for each subdomain, which are stitched together across boundaries and trained in parallel.
 Another class of domain decomposition techniques for PINNs consists of time marching methods, where, in general, a network is initially trained for a small subdomain, and then the subdomain is slowly expanded until it contains the entire domain.
-Methods in this class include backward-compatible PINNs, which consist of a single network shared by all subdomains [35].
-Also similar is the time marching approach in [36].
-Finally, one particularly successful technique is finite basis PINNs (FBPINNs) [37, 42, 43, 44], which use partition of unity functions to weight the output of neural networks in each domain.
-In comparison to other techniques, FBPINNs offer simplicity because no boundary conditions are needed between adjacent domains.
+Methods in this class include backward-compatible PINNs, which consist of a single network shared by all subdomains [mattey2022novel].
+Also similar is the time marching approach in [wight2020solving].
+Finally, a particularly successful technique is finite basis PINNs (FBPINNs) [moseley2023finite, dolean_finite_2024, dolean_multilevel_2024, heinlein2024multifidelity], which use partition of unity functions to weight the output of neural networks in each domain; these have also been extended to randomized neural networks [shang2025overlapping, anderson_elm-fbpinn_2024].
+In comparison to other techniques, FBPINNs offer simplicity because the boundary conditions between adjacent domains do not have to be enforced via additional penalty terms in the loss function.
 This work explores whether a similar technique can be used to enhance the training of a KAN for challenging cases.
-In addition to domain decomposition, many methods have been developed to improve the training of PINNs, including adaptive weighting schemes [45, 32], residual-based attention [46], adaptive residual point selection [47, 48, 49, 50, 51, 52], causality techniques [53], multifidelity PINNs [54, 55], and hierarchical methods to learn more accurate solutions progressively [56, 57, 58, 59, 60, 61, 62].
-A substantial advantage of the methods presented in this work is that these additional techniques could be applied in combination with FBKANs to improve the training further.
-In this paper, we first introduce finite basis KANs (FBKANs) in section 2.
-In sections 3 and 4 we highlight some of the features of FBKANs applied to data-driven problems and physics-informed problems, respectively.
+In addition to domain decomposition, many methods have been developed to improve the training of PINNs, including adaptive weighting schemes [mcclenny2020self, wang2022and], residual-based attention [anagnostopoulos2024residual], adaptive residual point selection [wu2023comprehensive, mao2023physics, hou2023enhancing, lau2024pinnacle, nabian2021efficient, gao2023failure, visser_pacmann_2024], causality techniques [wang2022causality], multifidelity PINNs [penwarden2022multifidelity, meng2019multifidelity], and hierarchical methods to learn more accurate solutions progressively [howard2023continual, howard2023stacked, wang2024multi, ainsworth2021galerkin, ainsworth2022galerkin, trask2022hierarchical, aldirany2024multi].
+A substantial advantage of the methodology presented in this work is that it can easily be combined with the aforementioned techniques to improve the training further.
+This paper is structured as follows: We first introduce finite basis KANs (FBKANs) in section˜2.
+In sections˜3 and 4, we highlight some of the features of FBKANs applied to data-driven and physics-informed problems, respectively.
 We show that FBKANs can increase accuracy over KANs.
 One important feature of FBKANs is that the finite basis architecture serves as a wrapper around a KAN architecture.
-While we have chosen to focus on KANs as described in [4], most available extensions of KANs could be considered instead to increase accuracy or robustness further.
-2 Methods
-2.1 KANs
-In [4], the authors proposed approximating a multivariate function $f(\mathbf{x})$ by a model of the form
+While we have chosen to focus on KANs as described in [liu2024kan], most available extensions of KANs could be considered instead to increase accuracy or robustness further.
+2 Methodology
+2.1 Kolmogorov-Arnold Networks
+In [liu2024kan], the authors proposed approximating a multivariate function $f(\mathbf{x})$ by a model of the form
 $f(\mathbf{x})\approx\sum_{i_{n_{l}-1}=1}^{m_{n_{l}-1}}\varphi_{n_{l}-1,i_{n_{l}},i_{n_{l}-1}}\left(\sum_{i_{n_{l}-2}=1}^{m_{n_{l}-2}}\ldots\left(\sum_{i_{2}=1}^{m_{2}}\varphi_{2,i_{3},i_{2}}\left(\sum_{i_{1}=1}^{m_{1}}\varphi_{1,i_{2},i_{1}}\left(\sum_{i_{0}=1}^{m_{0}}\varphi_{0,i_{1},i_{0}}(x_{i_{0}})\right)\right)\right)\ldots\right),$ (1)
-denoted as a Kolmogorov-Arnold network.
-Here, $n_{l}$ is the number of layers in the KAN, $\{m_{j}\}_{j=0}^{n_{l}}$ is the number of nodes per layer, and $\phi_{i,j,k}$ are the univariant activation functions.
-We denote the right-hand side of eq.
-1 as $\mathcal{K}(x)$.
-The activation functions are polynomials of degree $k$ on a grid with $g$ grid points.
+which they denote as a Kolmogorov-Arnold network.
+Here, $n_{l}$ is the number of layers in the KAN, $\{m_{j}\}_{j=0}^{n_{l}}$ is the number of nodes per layer, $\varphi_{i,j,k}$ are the univariate activation functions, and $\mathbf{x}=(x_{1},\ldots,x_{m_{0}})$.
+We denote the right-hand side of eq.˜1 as $\mathcal{K}(\mathbf{x})$.
+The activation functions $\varphi_{k-1,i_{k},i_{k-1}}$ in the $k$th layer are polynomials of degree $k$ on a grid with $g$ grid points.
 They are represented by a weighted combination of a basis function $b(x)$ and a B-spline,
-$\phi(x)=w_{b}b(x)+w_{s}\text{spline}(x)$ (2)
+$\varphi(x)=w_{b}b(x)+w_{s}\text{spline}(x),$ (2)
 where
@@
 $\text{spline}(x)=\sum_{i}c_{i}B_{i}(x).$
-Here, $B_{i}(x)$ is a polynomial of degree $k$. $c_{i},w_{b},$ and $w_{s}$ are trainable parameters.
+Here, $B_{i}(x)$ is a polynomial of degree $k$, and $c_{i}$, $w_{b}$, and $w_{s}$ are trainable parameters.
 KANs evaluate the B-splines on a precomputed grid.
-In one dimension, a domain $[a,b]$ a grid with $g_{1}$ intervals has grid points $\{t_{0}=a,t_{1},t_{2},\ldots,t_{g_{1}}=b\}$; cf.
-[4].
-Grid extension [4] allows for fitting a new, fine-grained spline to a coarse-grained spline, increasing the expressivity of the KAN.
-The coarse splines are transferred to the fine splines following the procedure in [4].
-In this work, we do not consider the method for enforcing sparsity as outlined in [4] and instead only consider the mean squared error in the loss function.
+In one dimension, in a domain $[a,b]$, a grid with $g_{1}$ intervals has grid points $\{t_{0}=a,t_{1},t_{2},\ldots,t_{g_{1}}=b\}$; cf.
+[liu2024kan].
+Grid extension [liu2024kan] allows for fitting a new, fine-grained spline to a coarse-grained spline, increasing the expressivity of the KAN.
+The coarse splines are transferred to the fine splines following the procedure described in [liu2024kan] and [rigas2024adaptive].
+In this work, we do not consider the method for enforcing sparsity of the trainable activation functions as outlined in [liu2024kan] and instead only consider the mean squared error (MSE) in the loss function.
 Many variations of KANs have been proposed recently.
-However, in this work, we consider only the formulation as outlined in [4], although we note that the domain decomposition method presented here would work with many variants of KANs.
+However, in this work, we consider only the formulation as outlined in [liu2024kan], although we note that the domain decomposition method presented here could be combined with many variants of KANs.
 2.2 Physics-informed neural networks
-A PINN is a neural network (NN) that is trained to approximate the solution of an initial-boundary value problem (IBVP) by optimizing loss terms accounting for initial conditions, boundary conditions, and a residual term using backpropagation to calculate derivatives.
-In particular, the residual term represents how well the output satisfies the governing differential equation [19].
+A physics-informed neural network (PINN) is a neural network (NN) that is trained to approximate the solution of an initial-boundary value problem (IBVP) by optimizing loss terms accounting for initial conditions, boundary conditions, and a residual term using backpropagation to calculate derivatives.
+In particular, the residual term represents how well the output satisfies the governing differential equation [raissi2019physics]; for early approaches of combining differential equations with neural networks, we refer to [dissanayake_neural-network-based_1994, lagaris_artificial_1998].
 More specifically, we aim at approximating the solution $f$ of a generic IBVP
-$f_{t}+\mathcal{N}_{x}[f]$ $=0,$ $x\in\Omega,t\in[0,T],$ (3)
-$f(x,t)$ $=g(x,t),$ $x\in\partial\Omega,t\in[0,T],$
-$f(x,0)$ $=u(x),$ $x\in\Omega,$
+$f_{t}+\mathcal{N}_{\mathbf{x}}[f]$ $=0,$ $x\in\Omega,t\in[0,T],$ (3)
+$f(\mathbf{x},t)$ $=g(\mathbf{x},t),$ $x\in\partial\Omega,t\in[0,T],$
+$f(\mathbf{x},0)$ $=u(\mathbf{x}),$ $x\in\Omega,$
 over the open domain $\Omega\in\mathbb{R}^{d}$ with boundary $\partial\Omega.$ Here, $x$ and $t$ represent the spatial and temporal coordinates, respectively, $\mathcal{N}_{x}$ is the differential operator with respect to $x$, and $g$ and $u$ are given functions representing the boundary and initial conditions, respectively.
-The PINN model is then trained to minimize the mean squared error for the initial conditions, boundary conditions, and residual (physics) sampled at $N_{ic},N_{bc},N_{r}$ data points.
+The PINN model is then trained to minimize the MSE for the initial conditions, boundary conditions, and residual (physics) sampled at $N_{ic},N_{bc},N_{r}$ data points.
 We denote the sampling points for the initial conditions, boundary conditions, and residual as $\{x_{ic}^{i},u(x_{ic}^{i})\}_{i=1}^{N_{ic}}$, $\{(x_{bc}^{i},t_{bc}^{i}),g(x_{bc}^{i},t_{bc}^{i})\}_{i=1}^{N_{bc}}$, and $\{(x_{r}^{i},t_{r}^{i})\}_{i=1}^{N_{r}}$, respectively.
-We then optimize the following weighted loss function with respect to the trainable parameters $\theta$ of the NN model $f_{\theta}$:
-$\mathcal{L}(\theta)$ $=\lambda_{ic}\mathcal{L}_{ic}(\theta)+\lambda_{bc}\mathcal{L}_{bc}(\theta)+\lambda_{r}\mathcal{L}_{r}(\theta),$ (4)
-$\mathcal{L}_{ic}(\theta)$ $=\frac{1}{N_{ic}}\sum_{i=1}^{N_{ic}}\left(f_{\theta}(x_{ic}^{i},0)-u(x_{ic}^{i})\right)^{2},$ (5)
-$\mathcal{L}_{bc}(\theta)$ $=\frac{1}{N_{bc}}\sum_{i=1}^{N_{bc}}\left(f_{\theta}(x_{bc}^{i},t_{bc}^{i})-g(x_{bc}^{i},t_{bc}^{i})\right)^{2},$ (6)
-$\mathcal{L}_{r}(\theta)$ $=\frac{1}{N_{r}}\sum_{i=1}^{N_{r}}\left(\frac{\partial}{\partial t}f_{\theta}(x_{r}^{i},t_{r}^{i})+\mathcal{N}_{x}[f_{\theta}(x_{r}^{i},t_{r}^{i})]\right)^{2},$ (7)
+Then, we optimize the following weighted loss function with respect to the trainable parameters $\theta$ of the NN model $f_{\theta}$:
+$\mathcal{L}(\theta)$ $=\lambda_{ic}\mathcal{L}_{ic}(\theta)+\lambda_{bc}\mathcal{L}_{bc}(\theta)+\lambda_{r}\mathcal{L}_{r}(\theta),$
+$\mathcal{L}_{ic}(\theta)$ $=\frac{1}{N_{ic}}\sum_{i=1}^{N_{ic}}\left(f_{\theta}(\mathbf{x}_{ic}^{i},0)-u(\mathbf{x}_{ic}^{i})\right)^{2},$
+$\mathcal{L}_{bc}(\theta)$ $=\frac{1}{N_{bc}}\sum_{i=1}^{N_{bc}}\left(f_{\theta}(\mathbf{x}_{bc}^{i},t_{bc}^{i})-g(\mathbf{x}_{bc}^{i},t_{bc}^{i})\right)^{2},$
+$\mathcal{L}_{r}(\theta)$ $=\frac{1}{N_{r}}\sum_{i=1}^{N_{r}}\left(\frac{\partial}{\partial t}f_{\theta}(\mathbf{x}_{r}^{i},t_{r}^{i})+\mathcal{N}_{\mathbf{x}}[f_{\theta}(\mathbf{x}_{r}^{i},t_{r}^{i})]\right)^{2},$
 where $\lambda_{ic},\lambda_{bc},\lambda_{r}$ are weights for the different loss terms.
 The choice of these weights can have a significant impact on the training process and model performance.
-While adaptive schemes are available for choosing the weights [45, 32, 63, 64], in this work, we set the weights manually.
-2.3 FBPINNs
-In FBPINNs [37, 42, 43, 44], we decompose the space domain $\Omega$ or the space-time domain $\Omega\times[0,T]$ into overlapping subdomains.
+While adaptive schemes are available for choosing the weights [mcclenny2020self, wang2022and, wang2022improved, howard2024conjugate], in this work, we set the weights manually to approximately scale the terms in the loss function to have equal magnitude.
+The values of the weights used are given in Appendix A.
+2.3 Finite basis physics-informed neural networks
+In finite basis physics-informed neural networks (FBPINNs) [moseley2023finite, dolean_finite_2024, dolean_multilevel_2024, heinlein2024multifidelity], we decompose the spatial domain $\Omega$ or the space-time domain $\Omega\times[0,T]$ into overlapping subdomains.
 Each overlapping subdomain $\Omega_{j}$ is the interior of the support of a corresponding function $\omega_{j}$, and all functions $\omega_{j}$ form a partition of unity.
 In particular, in the case of $L$ overlapping subdomains, we have
-$\Omega=\bigcup_{j=1}^{L}\Omega_{j},\quad{\rm supp}(\omega_{j})=\overline{\Omega_{j}},\text{ and}\quad\sum_{j=1}^{L}\omega_{j}\equiv 1\text{ in }\Omega.$
+[…]
 In one dimension, a uniform overlapping domain decomposition with overlap ratio $\delta>1$ is given by subdomains
 $\Omega_{j}=\left(\frac{(j-1)l-\delta l/2}{L-1},\frac{(j-1)l+\delta l/2}{L-1}\right),$
-where $l=\max(x)-\min(x)$.
-There are multiple ways to define the partition of unity functions. Here, we construct them based on the expression
-$\omega_{j}=\frac{\hat{\omega}_{j}}{\sum_{j=1}^{L}\hat{\omega}_{j}},$ (8)
+where $l=\max(x)-\min(x)$. Then, there are multiple ways to define the partition of unity functions. Here, we construct them based on the expression
+$\omega_{j}=\frac{\hat{\omega}_{j}}{\sum_{j=1}^{L}\hat{\omega}_{j}},$
 where
-$\hat{\omega}_{j}(x)=\begin{cases}1&L=1,\\ \left[1+\cos\left(\pi(x-\mu_{j})/\sigma_{j}\right)\right]^{2}&L>1,\end{cases}$ (9)
+$\hat{\omega}_{j}(x)=\begin{cases}1&L=1,\\ \left[1+\cos\left(\pi(x-\mu_{j})/\sigma_{j}\right)\right]^{2}&L>1,\end{cases}$
 with $\mu_{j}=l(j-1)/(L-1)$ and $\sigma_{j}=(\delta l/2)/(L-1)$ representing the center and half-width of each subdomain, respectively.
-An example of the one-dimensional partition of unity functions with $L=4$ is depicted in fig.
-2.
-Figure 2: Example the partition of unity functions on the domain $\Omega=[0,2]$ with $L=4$ subdomains.
-In multiple dimensions, $\mathbf{x}\in\mathbb{R}^{d}$, we then obtain
-$\hat{\omega}_{j}(x)=\begin{cases}1&L=1,\\ \Pi_{i=1}^{d}\left[1+\cos\left(\pi(x_{i}-\mu_{ij})/\sigma_{ij}\right)\right]^{2}&L>1,\end{cases}$ (10)
-with $\mu_{ij}$ and $\sigma_{ij}$ representing the center and half-width of each subdomain $j$ in each direction $i$.
-An example of of a single partition of unity function with $d=2$ and $L=4$ is shown in fig.
-3.
+An example of the one-dimensional partition of unity functions with $L=4$ is depicted in fig.˜2.
+Figure 2: Example of the partition of unity functions on the domain $\Omega=[0,2]$ with $L=4$ subdomains.
+In multiple dimensions, $\mathbf{x}\in\mathbb{R}^{d}$, we then obtain the partition of unity functions as the tensor product of the one-dimensional functions:
+$\hat{\omega}_{j}(\mathbf{x})=\begin{cases}1&L=1,\\ \Pi_{i=1}^{d}\left[1+\cos\left(\pi(x_{i}-\mu_{ij})/\sigma_{ij}\right)\right]^{2}&L>1,\end{cases}$
+with $\mu_{ij}$ and $\sigma_{ij}$ representing the center and half-width of each subdomain $j$ in each dimension $i$.
+An example of an overlapping domain decomposition and of a single partition of unity function with $d=2$ and $L=4$ are shown in fig.˜3.
 Figure 3: (Left) Example domain decomposition on the domain $\Omega=[-1,1]\times[-1,1]$ with $L=4$ subdomains. (Right) One example partition of unity function $\omega_{11}(x,y).$
 Then, the FBPINN architecture reads
-$f_{\theta}(x)=\sum_{j=1}^{L}\omega_{j}(x)f_{j,\theta^{j}}(x),$
-where the $f_{j,\theta^{j}}$ is the neural network with parameters $\theta^{j}$ that corresponds to the subdomain $\Omega_{j}$; it is localized to $\Omega_{j}$ by multiplication with the partition of unity function $\omega_{j}$, which is zero outside of $\Omega_{j}$.
+$f_{\theta}(\mathbf{x})=\sum_{j=1}^{L}\omega_{j}(x)f_{j}(\mathbf{x};\theta^{j}),$ (4)
+where $f_{j}(\cdot;\theta^{j})$ is the neural network with parameters $\theta^{j}$ that corresponds to the subdomain $\Omega_{j}$; it is localized to $\Omega_{j}$ by multiplication with the partition of unity function $\omega_{j}$, which is zero outside of $\Omega_{j}$.
 The FBPINN model is trained in the same way as the PINN model, i.e., using initial, boundary, and residual loss terms; cf.
-section 2.2.
-Note that, in the original FBPINNs paper [37], hard enforcement of initial and boundary conditions is employed, such that only the residual loss term remains.
-Moreover, in [43], the approach has been extended to multilevel domain decompositions, and in [44], this domain decomposition approach has been applied to long-range time-dependent problems.
-2.4 FBKANs
-In FBKANs, we replace the NN used in FBPINNs with a KAN The function approximation in eq. 1 then becomes
-$f(\mathbf{x})$ $\approx\sum_{j=1}^{L}\omega_{j}(x)\left[\sum_{i_{n_{l}-1}=1}^{m_{n_{l}-1}}\varphi^{j}_{n_{l}-1,i_{n_{l}},i_{n_{l}-1}}\left(\sum_{i_{n_{l}-2}=1}^{m_{n_{l}-2}}\ldots\left(\sum_{i_{1}=1}^{m_{1}}\varphi^{j}_{1,i_{2},i_{1}}\left(\sum_{i_{0}=1}^{m_{0}}\varphi^{j}_{0,i_{1},i_{0}}(x_{i_{0}})\right)\right)\ldots\right)\right]$ (11)
-$=\sum_{j=1}^{L}\omega_{j}(x)\mathcal{K}^{j}(x;\theta^{j})$
-where $\mathcal{K}^{j}(x;\theta^{j})$ denotes the $j^{th}$ KAN with trainable parameters $\theta^{j}$ and $\omega_{j}$ denotes the corresponding finite basis partition of unity function introduced in section 2.3.
-FBKANs are trained to minimize the loss function
+section˜2.2.
+Note that, in the original FBPINNs paper [moseley2023finite], hard enforcement of initial and boundary conditions is employed, such that only the residual loss term remains.
+Moreover, in [dolean_multilevel_2024] the approach has been extended to multilevel domain decompositions, in [heinlein2024multifidelity] this domain decomposition approach has been applied to long-range time-dependent problems, and in [shang2025overlapping, anderson_elm-fbpinn_2024] it has been extended to randomized neural networks.
+2.4 Finite basis Kolmogorov–Arnold networks
+In finite basis Kolmogorov–Arnold networks (FBKANs), we replace the NN used in FBPINNs with a KAN. The function approximation in eq.˜1 then becomes
+$f(\mathbf{x})$ $\approx\sum_{j=1}^{L}\omega_{j}(x)\left[\sum_{i_{n_{l}-1}=1}^{m_{n_{l}-1}}\varphi^{j}_{n_{l}-1,i_{n_{l}},i_{n_{l}-1}}\left(\sum_{i_{n_{l}-2}=1}^{m_{n_{l}-2}}\ldots\left(\sum_{i_{1}=1}^{m_{1}}\varphi^{j}_{1,i_{2},i_{1}}\left(\sum_{i_{0}=1}^{m_{0}}\varphi^{j}_{0,i_{1},i_{0}}(x_{i_{0}})\right)\right)\ldots\right)\right]$
+$=\sum_{j=1}^{L}\omega_{j}(x)\mathcal{K}_{j}(\mathbf{x};\theta^{j}),$ (5)
+where $\mathcal{K}_{j}(\mathbf{x};\theta^{j})$ denotes the $j^{th}$ KAN with trainable parameters $\theta^{j}$ and $\omega_{j}$ denotes the corresponding finite basis partition of unity function introduced in section˜2.3.
+FBKANs are trained to minimize the PINN loss function
 $\mathcal{L}(\theta)=\lambda_{ic}\mathcal{L}_{ic}(\theta)+\lambda_{bc}\mathcal{L}_{bc}(\theta)+\lambda_{r}\mathcal{L}_{r}(\theta)+\lambda_{data}\mathcal{L}_{data}(\theta),$ (6)
 composed of
-$\mathcal{L}_{ic}(\theta)$ $=\frac{1}{N_{ic}}\sum_{i=1}^{N_{ic}}\left(\sum_{j=1}^{L}\omega_{j}(x)\mathcal{K}^{j}(x_{ic}^{i};\theta^{j})-u(x_{ic}^{i})\right)^{2},$
-$\mathcal{L}_{bc}(\theta)$ $=\frac{1}{N_{bc}}\sum_{i=1}^{N_{bc}}\left(\sum_{j=1}^{L}\omega_{j}(x)\mathcal{K}^{j}(x_{bc}^{i};\theta^{j})-g(x_{bc}^{i},t_{bc}^{i})\right)^{2},$
-$\mathcal{L}_{r}(\theta)$ $=\frac{1}{N_{r}}\sum_{i=1}^{N_{r}}\left(\frac{\partial}{\partial t}\left[\sum_{j=1}^{L}\omega_{j}(x)\mathcal{K}^{j}(x_{r}^{i};\theta^{j})\right]+\mathcal{N}_{x}\left[\sum_{j=1}^{L}\omega_{j}(x)\mathcal{K}^{j}(x_{r}^{i};\theta^{j})\right]\right)^{2},$
-$\mathcal{L}_{data}(\theta)$ $=\frac{1}{N_{data}}\sum_{i=1}^{N_{data}}\left(\sum_{j=1}^{L}\omega_{j}(x)\mathcal{K}^{j}(x_{data}^{i};\theta^{j})-f(x_{data}^{i})\right)^{2}.$
+$\mathcal{L}_{ic}(\theta)$ $=\frac{1}{N_{ic}}\sum_{i=1}^{N_{ic}}\left(\sum_{j=1}^{L}\omega_{j}(\mathbf{x})\mathcal{K}_{j}(\mathbf{x}_{ic}^{i};\theta^{j})-u(\mathbf{x}_{ic}^{i})\right)^{2},$
+$\mathcal{L}_{bc}(\theta)$ $=\frac{1}{N_{bc}}\sum_{i=1}^{N_{bc}}\left(\sum_{j=1}^{L}\omega_{j}(\mathbf{x})\mathcal{K}_{j}(\mathbf{x}_{bc}^{i};\theta^{j})-g(\mathbf{x}_{bc}^{i},t_{bc}^{i})\right)^{2},$
+$\mathcal{L}_{r}(\theta)$ $=\frac{1}{N_{r}}\sum_{i=1}^{N_{r}}\left(\frac{\partial}{\partial t}\left[\sum_{j=1}^{L}\omega_{j}(\mathbf{x})\mathcal{K}_{j}(\mathbf{x}_{r}^{i};\theta^{j})\right]+\mathcal{N}_{\mathbf{x}}\left[\sum_{j=1}^{L}\omega_{j}(\mathbf{x})\mathcal{K}_{j}(\mathbf{x}_{r}^{i};\theta^{j})\right]\right)^{2},$
+$\mathcal{L}_{data}(\theta)$ $=\frac{1}{N_{data}}\sum_{i=1}^{N_{data}}\left(\sum_{j=1}^{L}\omega_{j}(\mathbf{x})\mathcal{K}_{j}(\mathbf{x}_{data}^{i};\theta^{j})-f(\mathbf{x}_{data}^{i})\right)^{2},$
+using the FBKAN model eq.˜5.
 Here, $\theta=\{\theta^{j}\}_{j=1}^{L}$ is the set of trainable parameters, and the loss function $\mathcal{L}$ contains both a term for data-driven training $\mathcal{L}_{data}$ and terms for physics-informed training $\mathcal{L}_{ic}$, $\mathcal{L}_{bc}$, and $\mathcal{L}_{r}$.
 In this way, FBKANs are adaptable to given problem characteristics.
-For FBKANs, the grid of each local KAN $\mathcal{K}^{j}$ is set separately on the corresponding subdomain $\Omega^{j}$.
+For FBKANs, the grid of each local KAN $\mathcal{K}_{j}$ is defined separately on each subdomain $\Omega^{j}$.
 To do so, we densely sample points $x\in\Omega$ and compute the partition of unity function $\omega_{j}(x)$.
 We then take the boundaries of each subdomain based on the partition of unity functions as
-$a^{j}$ $=\min_{x\text{ s.t. }\omega^{j}(x)>0.0001}x,$
-$b^{j}$ $=\max_{x\text{ s.t. }\omega^{j}(x)>0.0001}x.$
-This gives a custom grid fit to each subdomain.
-We evaluate the models using the relative $\ell_{2}$ error, calculated by
-$\frac{||f(x)-\sum_{j=1}^{L}\omega_{j}(x)\mathcal{K}^{j}(x;\theta^{j})||_{2}}{||f(x)||_{2}}.$
-Our implementation is based on Jax [65], using the Jax-KAN package [66] for the KAN implementation.
-For simplicity, we keep the hyperparameters of each KAN on each subdomain fixed, including the width and depth, reducing the total number of hyperparameters.
-However, in general, the parameters for each individual KAN $\mathcal{K}^{j}$ could be chosen differently.
+$a^{j}$ $=\min_{\begin{subarray}{c}x\text{ such that}\\ \omega^{j}(x)>10^{-4}\end{subarray}}x,$
+$b^{j}$ $=\max_{\begin{subarray}{c}x\text{ such that}\\ \omega^{j}(x)>10^{-4}\end{subarray}}x.$
+Hence, we obtain a custom grid for each subdomain.
+We evaluate the model performance using the relative $\ell_{2}$ error, calculated by
+[…]
+where $f$ is the ground truth solution.
+Our implementation is based on JAX [jax2018github], using the jaxKAN package [Rigas_jaxKAN_A_JAX-based_2024] for the KAN implementation.
+For simplicity, we employ the same hyperparameters for local KAN models defined on the subdomains, including the width and depth; this reduces the total number of hyperparameters.
+However, in general, the parameters for each individual KAN $\mathcal{K}_{j}$ could be chosen differently.
 3 Data-driven results
@@
 These results specifically illustrate how FBKANs can improve predictions for noisy data and multiscale oscillations.
-3.1 Test 1
-In this section, we consider a toy problem
+3.1 Data-driven test 1
+In this section, we consider a toy problem of fitting the function
 $f(x)=\exp[\sin(0.3\pi x^{2})],$ (7)
@@
 3.1.1 Scaling of FBKANs
-[…]
-Figure 4: Results with $N=2,8,$ and $32$ subdomains for eq.
-13.
+First, we consider clean data, $N_{data}=1\,200$, which is sampled uniformly from the interval $[0,8].$ We present results for training on $L=1,2,4,8,16,$ and $32$ subdomains.
+Note that the function described in eq.˜7 is highly oscillatory, which makes it difficult to capture its behavior with a small KAN.
+As shown in fig.˜4, increasing the number of subdomains from $L=2$ to $L=8$ and $L=32$ significantly decreases the pointwise error in the FBKAN predictions.
+Doubling the number of subdomains doubles the number of trainable parameters of the FBKAN model, increasing its expressivity.
+We observe approximately first order convergence of the relative $\ell_{2}$ error for a large number of subdomains in fig.˜4.
+Figure 4: Results with $L=2,8,$ and $32$ subdomains for eq.˜7.
 (a) Training data and plot of exact $f(x)$.
-(b) Loss curves (eq.
-12).
-(c) Plot of the outputs $f(x)$ and $\sum_{j=1}^{L}\omega_{j}(x)\mathcal{K}^{j}(x)$.
-(d) Pointwise errors $f(x)-\sum_{j=1}^{L}\omega_{j}(x)\mathcal{K}^{j}(x)$.
+(b) Loss curves (eq.˜6).
+(c) Plot of the outputs $f(x)$ and $\sum_{j=1}^{L}\omega_{j}(x)\mathcal{K}_{j}(x)$.
+(d) Pointwise errors $f(x)-\sum_{j=1}^{L}\omega_{j}(x)\mathcal{K}_{j}(x)$.
 (e) Scaling results for Test 1 with $L$ subdomains.
 3.1.2 FBKANs with noisy data
-One strength of KANs over MLPs is their increased accuracy for noisy data [4].
+One strength of KANs over MLPs is their increased accuracy for noisy data [liu2024kan].
 We now test FBKANs on noisy training data using four subdomains ($L=4$).
 Therefore, we sample $600$ training points from a uniform distribution in $[0,8]$ and evaluate $f(x)$.
-Then, we add Gaussian white noise with zero mean and varying magnitude to $f(x)$, up to an average relative magnitude of $20\%$.
-This becomes the training set.
+Then, we add Gaussian white noise with zero mean and varying magnitude to $f(x)$, up to an average relative magnitude of $20\,\%$, as the training data set.
 The method is then tested on $1000$ evenly spaced points in $[0,8]$ without noise.
-The training set and results are shown in fig.
-5.
-In the noisiest case, with relative noise of 18.1% added to the training set, the KAN yields a relative $\ell_{2}$ error of 0.1404, whereas the FBKAN yields a relative error of 0.0646.
-For all noise levels tested, the FBKAN consistently has a lower relative error than the plain KAN in fig.
-5, and the predictions are robust to noisy data.
+The training set and results are shown in fig.˜5.
+In the case with the largest noise, with relative noise of 18.1% added to the training set, the KAN yields a relative $\ell_{2}$ error of 0.1404, whereas the FBKAN yields a relative error of 0.0646.
+For all noise levels tested, the FBKAN consistently has a lower relative error than the plain KAN in fig.˜5, and the predictions are robust to noisy data.
 Figure 5: Results for Test 1 with noisy training data; cf.
-eq.
-13.
+eq.˜7.
 (a) Example training data and plot of exact $f(x)$ with $9.6\%$ mean relative noise.
-(b) Loss curves (eq.
-12) for an example training with $9.6\%$ mean relative noise.
-(c) Plot of the outputs $f(x)$ and $\sum_{j=1}^{L}\omega_{j}(x)\mathcal{K}^{j}(x)$ with $9.6\%$ mean relative noise.
-(d) Pointwise errors $f(x)-\sum_{j=1}^{L}\omega_{j}(x)\mathcal{K}^{j}(x)$ with $9.6\%$ mean relative noise.
+(b) Loss curves (eq.˜6) for an example training with $9.6\%$ mean relative noise.
+(c) Plot of the outputs $f(x)$ and $\sum_{j=1}^{L}\omega_{j}(x)\mathcal{K}_{j}(x)$ with $9.6\%$ mean relative noise.
+(d) Pointwise errors $f(x)-\sum_{j=1}^{L}\omega_{j}(x)\mathcal{K}_{j}(x)$ with $9.6\%$ mean relative noise.
 (e) Relative $\ell_{2}$ error of the KANs and FBKANs with respect to the magnitude of the noise added to the training data.
-3.2 Test 2
-We now consider the following equation
+3.2 Data-driven test 2
+Next, we consider fitting the function
 $f(x,y)=\sin(6\pi x^{2})\sin(8\pi y^{2}),$ (8)
@@
 The FBKAN has $L=4$ subdomains.
-The training set is composed of 10 000 points randomly sampled from a uniform distribution.
+The training set is composed of 10 00010\,000 points randomly sampled from a uniform distribution.
 In both cases, we begin training with $g=5$ grid points.
-In the grid extension case, the grid increases every 600 iterations as listed in table 4, and the learning rate drops by $20\%$ each time the grid is updated.
-As can be observed in fig.
-6, the final training loss for FBKAN-1 with a fixed grid is approximately the same as the training loss for KAN-2 with the grid extension approach, even though each FBKAN network has six times fewer grid points.
-This is also reflected in the relative errors reported in table 1.
-Comparing the KAN-1 and FBKAN-1 models in fig.
-7, KAN-1 struggles to capture the fine-scale features of the data accuratelyand has a larger relative error.
-As can be seen in fig.
-8, KAN-2 and FBKAN-2 both outperform their counterparts with a static grid, but FBKAN-2 is better able to capture the data than KAN-2.
-Figure 6: Training loss curves for Test 2.
+In the grid extension case, the grid increases every $600$ iterations as listed in table˜5, and the learning rate drops by $20\,\%$ each time the grid is updated.
+As can be observed in fig.˜6, the final training loss for FBKAN-1 with a fixed grid is approximately the same as the training loss for KAN-2 with the grid extension approach, even though each FBKAN network has six times fewer grid points.
+This is also reflected in the relative errors reported in table˜1.
+Comparing the KAN-1 and FBKAN-1 models in fig.˜7, KAN-1 struggles to capture the fine-scale features of the data accurately and has a larger relative error.
+As can be seen in fig.˜8, KAN-2 and FBKAN-2 both outperform their counterparts with a static grid, but FBKAN-2 is better able to capture the data than KAN-2.
+Figure 6: Training loss curves for the data-driven test 2.
 name grid type relative error
@@
 FBKAN-2 grid extension $2.27\times 10^{-2}$
-Table 1: Relative $\ell_{2}$ errors for physics-informed Test 2.
-Figure 7: Reference solution (left), predicted solutions (middle) and relative errors (right) for KAN-1 and FBKAN-1 for Test 2. Note that the error plots have different color scales.
-Figure 8: Reference solution (left), predicted solutions (middle) and relative errors (right) for KAN-2 and FBKAN-2 for Test 2. Note that the error plots have different color scales.
+Table 1: Relative $\ell_{2}$ errors for data-driven test 2.
+Figure 7: Reference solution (left), predicted solutions (middle) and relative errors (right) for KAN-1 and FBKAN-1 for data-driven test 2. Note that the error plots have different color scales.
+Figure 8: Reference solution (left), predicted solutions (middle) and relative errors (right) for KAN-2 and FBKAN-2 for data-driven test 2. Note that the error plots have different color scales.
 4 Physics-informed results
-4.1 Test 3
+4.1 Physics-informed test 1
 As the first physics-informed example, we consider the following one-dimensional problem with multiscale features:
@@
 $f(0)$ $=0,$
-on $x\in[-4,4]$ with exact solution
+for $x\in[-4,4]$. The analytical solution reads
 $f(x)=\sin(4x)+\sin(40x).$ (9)
 We sample $400$ collocation points from a uniform distribution ($N_{r}=400$) at every training iteration.
-The grid starts with $g=5$ and increases by $5$ every 1 000 iterations.
+We start with a grid with $g=5$ and extend it by $5$ every $1000$ iterations.
 The KAN reaches a relative $\ell_{2}$ error of 0.2407 with $g_{final}=20$, whereas the FBKAN for four subdomains ($L=4$) reaches a relative $\ell_{2}$ error of 0.0898 for $g_{final}=20$.
 For eight subdomains ($L=8$), the FBKAN reaches a relative error of 0.0369 with $g_{final}=15$ and 0.0662 with $g_{final}=10$.
-Figure 9: Results for Test 3 with $L=8$. The FBKAN clearly outperforms the KAN, reaching a lower loss with fewer grid points.
-4.2 Test 4
-We now test the Helmholtz equation
+Figure 9: Results for physics-informed test 1 with $L=8$. The FBKAN clearly outperforms the KAN, reaching a lower loss with fewer grid points.
+4.2 Physics-informed test 2
+Next, we test the Helmholtz equation
 $\frac{\partial^{2}f}{\partial y^{2}}+\frac{\partial^{2}f}{\partial x^{2}}+k_{h}^{2}f-q(x,y)$ $=0,\ (x,y)\in[-1,1]\times[-1,1],$
@@
 with
-$q(x,y)=-(a_{1}\pi)^{2}\sin(a_{1}\pi x)\sin(a_{2}\pi y)-(a_{2}\pi)^{2}\sin(a_{1}\pi x)\sin(a_{2}\pi y)+k_{h}\sin(a_{1}\pi x)\sin(a_{2}\pi y).$
+$q(x,y)=-(a_{1}\pi)^{2}\sin(a_{1}\pi x)\sin(a_{2}\pi y)-(a_{2}\pi)^{2}\sin(a_{1}\pi x)\sin(a_{2}\pi y)+k_{h}^{2}\sin(a_{1}\pi x)\sin(a_{2}\pi y)$
+which is obtained by using the solution
+$f(x,y)=\sin(a_{1}\pi x)\sin(a_{2}\pi y).$
 In our tests, we vary $a_{1}$ and $a_{2}$.
@@
 Additionally, we test a third case, denoted by KAN-3 / FBKAN-3, where we consider the same hyperparameters as KAN-1 / FBKAN-1 but with width 5 in the hidden layer.
-The relative $\ell_{2}$ errors are reported in table 2.
-In all cases, the FBKAN outperforms the corresponding KAN.
-The grid extension KAN cases, where we increase the grid size over the training, comes at the expense of increasing the computational time by a factor of 1.4-1.5 in our tests.
-For $a_{1}=1,a_{2}=4$, the smaller FBKAN-3 network is sufficient to represent the solution accurately.
+The relative $\ell_{2}$ errors are reported in table˜2.
+In all cases, the FBKAN outperforms the corresponding KAN model.
+The grid extension KAN cases, where we increase the grid size during training, comes at the expense of increasing the computational time by a factor of 1.41.51.4--1.5 in our tests.
+For $a_{1}=1$ and $a_{2}=4$, the smaller FBKAN-3 network is sufficient to represent the solution accurately.
 However, for larger values of $a_{1}$ and $a_{2}$, the larger network in FBKAN-1 is necessary.
-The results for KAN-1/FBKAN-1 are shown in figs.
-10 and 11 for $a_{1}=1,a_{2}=4$ and $a_{1}=4,a_{2}=4$, respectively.
-In fig.
-12, we consider $a_{1}=a_{2}=6$ with $L=4,9$ and $16$, demonstrating the further refinement possible with additional subdomains.
+The results for KAN-1 / FBKAN-1 are shown in figs.˜10 and 11 for $(a_{1},a_{2})=(1,4)$ and $(a_{1},a_{2})=(4,4)$, respectively.
+In fig.˜12, we consider $a_{1}=a_{2}=6$ with $L=4,9$ and $16$, demonstrating the further refinement possible with additional subdomains.
 $a_{1}=1$, $a_{2}=4$ $a_{1}=4$, $a_{2}=4$ $a_{1}=6$, $a_{2}=6$
@@
 FBKAN-3 0.0214 0.2760 0.9797
-Table 2: Relative $\ell_{2}$ errors for physics-informed Test 3.
-Figure 10: Results for Test 4 with KAN-1 and FBKAN-1, and $a_{1}=1,a_{2}=4$, $L=4$.
+Table 2: Relative $\ell_{2}$ errors for physics-informed test 2.
+Figure 10: Results for physics-informed test 2 with KAN-1 and FBKAN-1, and $a_{1}=1,a_{2}=4$, $L=4$.
 The solution along the line $x=0$ is given in the bottom left subfigure.
 Note that the error plots have different color scales.
-Figure 11: Results for Test 4 with KAN-1 and FBKAN-1, and $a_{1}=4,a_{2}=4$, $L=4$.
+Figure 11: Results for physics-informed test 2 with KAN-1 and FBKAN-1, and $a_{1}=4,a_{2}=4$, $L=4$.
 The solution along the line $x=0$ is given in the bottom left subfigure.
 Note that the error plots have different color scales.
-Figure 12: Results for Test 4 with KAN-1 and FBKAN-1 and $a_{1}=a_{2}=6$, with $L=4$ and $L=16$.The solution along the line $x=0$ is given in the middle left subfigure.
+Figure 12: Results for physics-informed test 2 with KAN-1 and FBKAN-1 and $a_{1}=a_{2}=6$, with $L=4$ and $L=16$.The solution along the line $x=0$ is given in the middle left subfigure.
 Note that the error plots have different color scales.
-4.3 Test 5
+4.3 Physics-informed test 3
 Finally, we consider the wave equation
@@
 $f(x,t)=\sin(\pi x)\cos(c\pi t)+0.5\sin(4\pi x)\cos(4c\pi t).$
-We first consider $c=\sqrt{2}$.
-The KAN has a relative $\ell_{2}$ error of 0.1402, and the FBKAN with $L=4$ has a relative $\ell_{2}$ error of 0.0153, as illustrated in fig.
-13.
-We then consider the harder case with $c=2$, shown in fig.
-14.
+For $c=\sqrt{2}$, the KAN model has a relative $\ell_{2}$ error of 0.1402, and the FBKAN with $L=4$ has a relative $\ell_{2}$ error of 0.0153, as illustrated in fig.˜13.
+We then consider the more challenging case with $c=2$, shown in fig.˜14.
 The KAN has a relative $\ell_{2}$ error of 0.1778 and the FBKAN with $L=4$ has a relative $\ell_{2}$ error of 0.0587.
-Figure 13: Results for Test 5 with $c=\sqrt{2}$ and $L=4$. Note that the error plots have different color scales.
-Figure 14: Results for Test 5 with $c=2$ and $L=4.$ Note that the error plots have different color scales.
+Figure 13: Results for physics-informed test 3 with $c=\sqrt{2}$ and $L=4$. Note that the error plots have different color scales.
+Figure 14: Results for physics-informed test 3 with $c=2$ and $L=4.$ Note that the error plots have different color scales.
+5 Multilevel FBKANs
+Multilevel FBPINNs (MLFBPINNs) were introduced in [dolean_multilevel_2024] to extend FBPINNs to include multiple levels of domain decomposition simultaneously.
+The motivation is to allow for approximating components on various different scales.
+Notably, it had been observed that low-frequency components, also denoted coarse components, cannot be learned well with the one-level approach if the numbers of subdomains are being increased; this makes it necessary to add one or multiple coarser corrections; cf.
+[dolean_finite_2024, dolean_multilevel_2024].
+In MLFBPINNs, the FBPINN architecture in eq.˜4 is modified to accommodate $N$ levels of domain decomposition with each level $l$ having an overlapping domain decomposition with $J^{(l)}$ subdomains
+$D^{(l)}=\left\{\Omega_{j}^{(l)}\right\}_{j=1}^{J^{(l)}}.$
+We take partition of unity functions $\omega_{j}^{(l)}$. As in the single level case, we then require
+[…]
+The MLFBPINN architecture is given by
+$f_{\theta}(\mathbf{x})=\frac{1}{N}\sum_{l=1}^{N}\sum_{j=1}^{J^{(l)}}\omega_{j}^{(l)}(x)f_{j}^{(l)}(\mathbf{x};\theta_{j}^{(l)}),$ (10)
+where again $f_{j}^{(l)}(\cdot;\theta_{j}^{(l)})$ is the neural network with parameters $\theta_{j}^{(l)}$ that corresponds to the subdomain $\Omega_{j}^{(l)}$ in level $l$.
+The extension to KANs, resulting in MLFBKANs, is then straightforward:
+$f(\mathbf{x})\approx\frac{1}{N}\sum_{l=1}^{N}\sum_{j=1}^{J^{(l)}}\omega_{j}^{(l)}(x)\mathcal{K}_{j}^{(l)}(\mathbf{x};\theta^{j}),$
+where $\mathcal{K}_{j}^{(l)}$ is the KAN corresponding to the subdomain $\Omega_{j}^{(l)}$ on level $l$.
+In this section, we show that MLFBKANs can improve training over FBKANs for solutions with higher or multiple frequencies by considering the Helmholtz equation from section˜4.2 and a two-dimensional multiscale Laplacian problem inspired by [dolean_multilevel_2024].
+In order to indicate how many subdomains on the different levels are being used we employ the notation $L=n_{1},n_{2},\ldots$; it means that $n_{1}$ subdomains are used on the first level, $n_{2}$ on the second level, and so on.
+5.1 Multilevel physics-informed test 1
+We first consider the Helmholtz equation from section˜4.2, choosing larger values of $a_{1}=a_{2}$ to increase the computational intensity of training the MLFBKAN.
+We consider $a_{1}=a_{2}=8$ and $a_{1}=a_{2}=10$.
+Error results are reported in table˜3.
+We compare the results across different architectures in figs.˜15 and 16.
+In both cases, the MLFBKAN with three or four levels outperforms the single level FBKAN, even with 36 subdomains.
+Indeed, for $a_{1}=a_{2}=8$, even the MLFBKAN with two levels outperforms the single level FBKANs.
+These results suggest that MLFBKANs can be more accurate than FBKANs for highly oscillatory problems, while reducing the total number of subdomains needed for accurate results.
+Figure 15: Results for multilevel physics-informed test 1 with $a_{1}=1,a_{2}=8$. Note that the error plots have different color scales.
+Figure 16: Results for multilevel physics-informed test 1 with $a_{1}=1,a_{2}=10$. Note that the error plots have different color scales.
+5.2 Multilevel physics-informed test 2
+Finally, we consider a multiscale Laplacian problem also employed in [dolean_multilevel_2024]. We take
+$-\nabla^{2}u(x,y)$ $=f(x,y),$ $(x,y)\in[0,1]\times[0,1],$
+$u(0,y)$ $=u(1,y)=0,$ $y\in[0,1],$
+$u(x,0)$ $=u(x,1)=0,$ $x\in[0,1],$
+with
+$f(x,y)=\frac{2}{M}\sum_{i=1}^{M}(2^{i}\pi)^{2}\sin(2^{i}\pi x)\sin(2^{i}\pi y).$
+The exact solution is then given by
+$u(x,y)=\frac{1}{M}\sum_{i=1}^{M}\sin(2^{i}\pi x)\sin(2^{i}\pi y).$
+From table˜3, for the case of $M=5$, the MLFBKANs offer a significant reduction of error over the FBKANs.
+These predictions and errors are shown in fig.˜17.
+The errors indicate that the solution can be accurately captured with the finest resolution domain decomposition having 16 subdomains.
+Figure 17: Results for multilevel physics-informed test 2 with $M=5$. Note that the error plots have different color scales.
+$a_{1}=a_{2}=8$ $a_{1}=a_{2}=10$ $M=5$
+KAN 0.89089 0.92729 1.08556
+FBKAN, L=4 0.91962 0.87100 0.33474
+FBKAN, L=16 0.28707 0.72251 0.04175
+FBKAN, L=36 0.33562 0.51766 0.09343
+MLFBKAN, N = 2, L=1, 4 0.24593 0.87294 0.10461
+MLFBKAN, N = 3, L=1, 4, 16 0.06353 0.27380 0.02926
+MLFBKAN, N = 4, L=1, 4, 16, 36 0.04231 0.12012 0.03066
+Table 3: Relative $\ell_{2}$ errors for the multilevel physics-informed test case 1 (middle columns) and test case 2 (right column).
 6 Conclusions
-We have developed domain decomposition-based KAN models for data-driven and physics-informed training with KANs; in accordance with FBPINNs, we denote them as FBKANs.
-FBKANs are scalable to complex problems and have a strong advantage over other domain decomposition-based approaches in that they do not require enforcement of transmission conditions between the subdomains via the loss function.
-They allow accurate training using an ensemble of small KANs combined using partition of unity functions, instead of a single large network.
-One advantage of FBKANs is that they can be combined with existing techniques to improve the training of KANs and PI-KANs, including residual-based attention weights as introduced in [7], cKANs [13], deep operator KANs [6], and others.
-These methods would be directly compatible with the FBKAN framework.
+We have developed domain decomposition-based architectures for data-driven and physics-informed training with KANs; in accordance with finite basis physics-informed neural networks (FBPINNs), we denote them as finite basis Kolmogorov–Arnold networks (FBKANs).The finite basis approach enables scalability for complex problems, and they have a strong advantage over other domain decomposition-based approaches in that they do not require enforcement of transmission conditions between the subdomains via the loss function.
+They allow accurate training using an ensemble of small KANs combined using partition of unity functions, instead of a single large KAN model.
+We also considered MLFBKANs, following the multilevel FBPINNs [dolean_multilevel_2024] approach.
+Compared to the one-level approach, MLFBKANs show improvements for training with larger numbers of subdomains, which is particularly interesting for high or multi frequency problems.
+Furthermore, FBKANs can be combined with existing techniques to improve the training of KANs and physics-informed KANS (PI-KANs), including residual-based attention weights as introduced in [shukla2024comprehensive], cKANs [ss2024chebyshev], deep operator KANs [abueidda2025deepokan], and others.
 In future work, we will further examine the scalability of FBKANs and consider their application to higher dimensional problems.
-We will also consider multilevel FBKANs, following the multilevel FBPINNs [43] approach.
-Multilevel FBPINNs show improvement over FBPINNs for a large number of subdomains by providing a mechanism for global communication between subdomains.
-Multilevel FBKANs could offer a similar advantage, allowing robust training with an increasing number of subdomains.
 7 Code and data availability
-All code, trained models, and data required to replicate the examples presented in this paper will be released upon publication.
-Meanwhile, we have released code and Google Colab tutorials for FBKANs in Neuromancer [67] at https://github.com/pnnl/neuromancer/tree/feature/fbkans/examples/KANs, for the reader to explore the ideas implemented in this work.
-7 Acknowledgements
-The KAN diagrams in fig.
-1 were developed with pykan [4].
+All code and data required to replicate the examples presented in this paper will be released upon publication.
+Meanwhile, we have released code and Google Colab tutorials for FBKANs in Neuromancer [Neuromancer2023] at https://github.com/pnnl/neuromancer/tree/feature/fbkans/examples/KANs for the reader to explore the ideas implemented in this work.
+8 Acknowledgments
+The KAN diagrams in fig.˜1 were developed with pykan [liu2024kan].
 This project was completed with support from the U.S.
@@
 Appendix A Training parameters
-All results in this paper are implemented in JAX [65] using the Jax-KAN [66] KAN implementation.
+All results in this paper are implemented in JAX [jax2018github] using the Jax-KAN [Rigas_jaxKAN_A_JAX-based_2024] KAN implementation.
 All networks are trained with the ADAM optimizer.
-For all FBKANs we take the domain overlap $\delta=1.9$.
-A.1 Test 1
-parameter section 3.1.1 section 3.1.2
+For all FBKANs we take the domain overlap $\delta=1.9$, based on the values used in [dolean_multilevel_2024].
+A.1 Data-driven test 1
+parameter section˜3.1.1 section˜3.1.2
 KAN architecture [1, 5, 1] [1, 5, 1]
-$L$ 4 - 32 4
+$L$ 4 – 32 4
 $g$ 5 5
@@
 learning rate 0.04 0.04
-Iterations 4000 4000
+# iterations $4000$ $4000$
 $N_{data}$ $1200$ $600$
-Table 3: Hyperparameters used for the results in section 3.1.
-A.2 Test 2
+Table 4: Hyperparameters used for the results of the data-driven test 1 in section˜3.1.
+A.2 Data-driven test 2
 parameter KAN-1 / FBKAN-1 KAN-2 / FBKAN-2
@@
 $g$ 5 [5, 10, 25, 30]
-$g$ schedule - [0, 600, 1200, 1800]
+$g$ schedule - [0, $600$, $1200$, $1800$]
 $k$ 3 3
@@
 learning rate scale - 0.8
-Iterations 2400 2400
-$N_{data}$ 10000 10000
-Table 4: Hyperparameters used for the results in section 3.2.
+# iterations $2400$ $2400$
+$N_{data}$ 10 00010\,000 10 00010\,000
+Table 5: Hyperparameters used for the results of the data-driven test 2 in section˜3.2.
 The grid ($g$) schedule denotes the iterations at which the grid is updated.
@@
 KAN-1 / FBKAN-1 use fixed grids.
-A.3 Test 3
+A.3 Physics-informed test 1
 parameter
@@
 $g$ [5, 10, 15, 20]
-$g$ schedule [0, 1000, 2000, 3000]
+$g$ schedule [0, $1000$, $2000$, $3000$]
 $k$ 3
@@
 learning rate scale 0.8
-Iterations 4000
+# iterations $4000$
 $N_{r}$ $400$
@@
 $\lambda_{ic}$ 1
-Table 5: Hyperparameters used for the results in section 4.1.
+Table 6: Hyperparameters used for the results of the physics-informed test 1 in section˜4.1.
 The grid ($g$) schedule denotes the iterations at which the grid is updated.
 The learning rate scale denotes the change to the learning rate at each grid update.
-A.4 Test 4
+A.4 Physics-informed test 2
 parameter KAN-1 / FBKAN-1 KAN-2 / FBKAN-2 KAN-3 / FBKAN-3
 KAN architecture [2, 10, 1] [2, 10, 1] [2, 5, 1]
-$L$ 4-16 4 4
+$L$ 4 – 16 4 4
 $g$ 5 [5, 10, 15] 5
-$g$ schedule - [0, 3000, 6000] -
+$g$ schedule - [0, $3000$, $6000$] -
 $k$ 5 3 5
@@
 learning rate scale - 0.8 -
-Iterations $a_{1}=1,a_{2}=4$: 10000 $a_{1}=4,a_{2}=4$: 10000 $a_{1}=6,a_{2}=6$: 30000 10000 10000
+# iterations
+$a_{1}=1,a_{2}=4$: 10 00010\,000
+$a_{1}=4,a_{2}=4$: 10 00010\,000
+$a_{1}=6,a_{2}=6$: 30 00030\,000
+10 00010\,000 10 00010\,000
 $N_{r}$ 800 800 800
@@
 $\lambda_{r}$ 0.01 0.01 0.01
-$\lambda_{ic}$ 1 1 1
-Table 6: Hyperparameters used for the results in section 4.2.
+$\lambda_{bc}$ 1 1 1
+Table 7: Hyperparameters used for the results of the physics-informed test 2 in section˜4.2.
 The grid ($g$) schedule denotes the iterations at which the grid is updated.
 The learning rate scale denotes the change to the learning rate at each grid update.
-A.5 Test 5
+A.5 Physics-informed test 3
 parameter $c=\sqrt{2}$ $c=2$
@@
 initial learning rate 0.001 0.0005
-Iterations 60000 120000
+# iterations 60 00060\,000 120 000120\,000
 $N_{r}$ $1000$ $1200$
@@
 $\lambda_{bc}$ 1 1
-Table 7: Hyperparameters used for the results in section 4.3.
+Table 8: Hyperparameters used for the results of the physics-informed test 3 in section˜4.3.
+A.6 Multilevel physics-informed test 1
+parameter
+KAN architecture [2, 10, 1]
+$L$ 4 – 36
+NN 1 – 4
+$g$ 5
+$g$ schedule -
+$k$ 5
+initial learning rate 0.005
+learning rate scale -
+# iterations 30 00030\,000
+$N_{r}$ 800
+$N_{bc}$ 400
+$\lambda_{r}$ 0.01
+$\lambda_{bc}$ 1
+Table 9: Hyperparameters used for the results of the multilevel physics-informed test 1 in section˜5.1.
+The grid ($g$) schedule denotes the iterations at which the grid is updated.
+The learning rate scale denotes the change to the learning rate at each grid update.
+A.7 Multilevel physics-informed test 2
+parameter
+KAN architecture [2, 10, 1]
+$L$ 4 – 36
+NN 1 – 4
+$g$ 5
+$g$ schedule -
+$k$ 5
+initial learning rate 0.005
+learning rate scale -
+# iterations 30 00030\,000
+$N_{r}$ 800
+$N_{bc}$ 400
+$\lambda_{r}$ 0.001
+$\lambda_{bc}$ 1
+Table 10: Hyperparameters used for the results of the multilevel physics-informed test 2 in section˜5.2.
+The grid ($g$) schedule denotes the iterations at which the grid is updated.
+The learning rate scale denotes the change to the learning rate at each grid update.